Proof the the equation in the black writint using the equation in blue writing
ectio P(AUB) P(A)+P(B)-P(Ane). AUB P(AUB) P(A) +P(B) AAR
Answer #1

If we express the event AUB as the union of two exclusive events A and A^{^{c}}cap B .As is clear from the figure.

ADB Anß

AUB = A cup (A^{c ca ca} cap B)

Since the events on right hand side are exclusive, by axiom. 3 (i.e. P(AUB) = P(A) + P(B) if A and B are mutually exclusive events)

P(A cup B) = P(A) + P ( A^{c} cap B) ----------(1)

We know that,

P ( A^{c} cap B) = P (B) - P( A cap B) ------------(2)

Putting the value of P ( A^{c} cap B) from (2) in (1), we get

P(A U B) = P(A) + P(B) _ P(Αλ B )